1) Certain elements emit light of a specific color when they are burned. When an potassium solution is burned in a aflame test, the energy of the light emitted is 4.909x10^-19 J. Calculate the wavelength in (nm) for this light.

2) a. A baseball with a mass 0.143 kg is thrown towards a batter at a velocity of 42.5 m/s, calculate the wavelength (in m) associated with the baseball's motion.

b. How does the λ (lambda) compare in size to the baseball ( diameter = 0.08 m)?

3) a. Calculate the wavelength (in m) associated with an electron traveling at the same velocity, 42.5 m/s. ( The mass of the electron is 9.1095 x 10^-31 kg)

b. How does the λ compare in size to the electron ( Diameter = 1x10^-10 m) ?
Answer #1

1)

use:

E = h*c/lambda

4.909*10^-19J =(6.626*10^-34 J.s)*(3.0*10^8 m/s)/lambda

lambda = 4.05*10^-7 m

= 405 nm

Answer: 405 nm

2)

a)

so, we have:

v = 42.5 m/s

mass, m = 0.143 kg

use:

lambda = h / (m*v)

lambda = (6.626*10^-34) / (0.143*42.5)

lambda = 1.09*10^-34 m

Answer: 1.09*10^-34 m

  

b)

ratio = 0.08 / 1.09*10^-34

= 7.4*10^32

base ball is 7.4*10^32 time larger

Only 1 question at a time please